General Meeting Summary 10/4/17

From cua_phy
Jump to navigation Jump to search

SHMS AEROGEL CHERENKOV

  • Overall, detector with SP-30 tray shows same number of photoelectrons (13-15) as before and is working fine
  • Discussion about noisy channels
  • changing HV card seems to have fixed the problem
  • also swapped PMT bases and changed HV cables
  • Discussion about additional peak in photoelectron spectrum
  • Issue fixed by fixing the Analyzer - problem was decoding of fADC
  • Action items:
  • take long cosmic run to verify that all is working
  • Change channel 6 PMT with one of the spares that were at CERN


PBWO4 SICCAS 2017

  • Discussion about best tool for measuring dimensions of crystals
  • best seems digital height gage by Mitutoyo


DIRC

  • Making good progress with preparations for bar movement from SLAC to JLab - currently planned for November


CERN DETECTORS

  • Main efforts currently to get lab at CUA setup
  • found issue with gantry not fitting into elevator to get it from basement to second floor
  • possible solutions include to lift gantry to second floor from outside building, modifying elevator doors, etc. - discussions with moving/rigging company ongoing
  • vacuum lines being installed
  • Discussion about lower luminosity running at LHC for rest of the year - high luminosity anticipated in 2018 (needed for pixel detector studies)


DATA ANALYSIS

  • DVCS deadtime analysis near completion
  • E12-09-011 rate projections with acctual energies in progress


NEXT MEETING: WED. 11 OCTOBER 2017 at 10:00 AM (EDT)